The coefficient of kinetic friction between a suit
Answer #1

mass of the suitcase,m=50 kg

Work done=620 J

coeffecient of kinetic frction,uk=0.272

Work done=Force*distance

Force needed is to as great as the friction force,F=uk*m*g=50*0.272*9.8=133.28 N

distance,d=E/F=620/133.28=4.65 m